# Exploring Representations and Interventions in Time Series Foundation Models

### v2

@@ 3.2 Identifying, Localizing and Steering Time Series Concepts
Once we have identified that a feature is linearly represented in the latent space of the $\mathcal{M}$, we can use steering interventions to manipulate the latent space and generate time series that reflect intended concepts. For instance, to introduce periodicity to a constant time series, we can utilize a steering matrix $\mathbf{S}$, as illustrated in Figure 2. By strategically intervening in $\mathcal{M}$ using this steering matrix, we can bias its outputs towards predicting periodic time series. To construct a steering matrix, we first derive steering vectors $\mathbf{S}_{i}\in\mathbb{R}^{N\times D}$, for each layer $i$. These vectors represent the change that activations in layer $i$ must undergo such that $\mathcal{M}$ produces periodic outputs. $\mathbf{S}_{i}$ is simply the difference between the median activation matrix of the constant time series $\mathbf{M}_{i_{c}}$, from that of sinusoids $\mathbf{M}_{i_{s}}$. We stack these vectors for all layers to derive the steering matrix. This matrix allows us to simultaneously intervene across multiple tokens and layers during inference, which we found to be more effective than single-token interventions. During inference, to steer model predictions, at each layer $i$, we update its hidden representation as follows: $\mathbf{h}_{i}\leftarrow\mathbf{h}_{i}+\alpha\mathbf{S}_{i},$ where $\alpha\in\mathbb{R}$ is a scalar that controls the strength of the intervention.
Research Questions and Experimental Setup.
Through our experiments, we aim to answer the following research questions: (RQ4) Do TSFMs represent concepts associated with specific data-generating functions distinctly in the latent space? (RQ5) Are these learned concepts localized to specific layers and tokens within TSFMs? (RQ6) Can we leverage these learned concepts to bias model predictions towards intended outcomes? For example, can we add periodicity or an upward trend to a constant time series? (RQ7) Is it possible to combine multiple steering interventions to manipulate model predictions towards complex compositions of various concepts? For instance, can we steer a model to add both trend and periodicity to a constant signal?
To address these research questions, we will leverage the techniques outlined in previous sections. We also explore two alternative modalities of intervention: (1) deriving steering vectors using the mean of hidden activations rather than the median, and (2) steering a single token versus all tokens throughout the model.
Refer to caption Refer to caption Refer to caption Refer to caption
(i) Pattern (ii) Trend (iii) Periodicity (iv) Amplitude
Figure 4: These heatmaps visualize the linear separability of various concepts at the patch level. Linear separability refers to the Linear Discriminant Ratio (LDR) computed from model embedding statistics for each predicted class: constant versus sinusoidal patterns (i), increasing versus decreasing trends (ii), high versus low periodicity (iii), and high versus low sinusoidal amplitude (iv). Lighter shades indicate higher separability (dark blue $\rightarrow$ low LDR, yellow $\rightarrow$ high LDR). These heatmaps show that certain concepts represented by MOMENT-Large are linearly separable and that this separability is not consistent but rather emerges at specific layers in the model.
Analyzed Time Series Foundation Models.
@@
While our methods are broadly applicable to a wide range of transformer-based foundation models, we focus on two prominent TSFM families for brevity: MOMENT [^1] https://github.com/moment-timeseries-foundation-model/moment (Goswami et al., 2024) and Chronos [^2] https://github.com/amazon-science/chronos-forecasting (Ansari et al., 2024). Both these models are fully open-source, come in different sizes, yet have fundamentally different design choices. For example, Chronos is forecasting on encoder-decoder transformer model which takes discretized time series as input, whereas MOMENT is a multi-task, encoder-only model which takes continuous time series patches as input. Since only the Large variant of MOMENT is publicly available at the time of writing this paper, we supplement our representation analysis results with Moirai (Woo et al., 2024) [^3] https://github.com/SalesforceAIResearch/uni2ts, another another popular TSFM which comes in different sizes. More information on model parameters can be found in Appendix C.
@@
(i) Trend, $\alpha=1.0$ (ii) Trend & periodicity, $\alpha=0.5$ (iii) Periodicity, $\alpha=0$
@@
Figure 5: Visualization of the MOMENT’s reconstruction and the Chronos’s forecasting predictions (bottom), with concept steering applied in the latent space (blue) and the baseline without concept steering (orange). Concept steering effectively transforms concepts in the latent space, resulting in model predictions that align with the intended concepts introduced. Steering results are illustrated for the following experiments: steering a constant signal input to produce (i) a sinusoidal output, (ii) a constant signal with an increasing(slope $>$ 0), and (iii) decreasing trend (slope $<$ 0). For compositional steering experiments, $\alpha$ controls the strength of sinusoidal and increasing trend concepts. When $\alpha=0.5$, models are steered towards a combination of increasing trend and sinusoidal pattern. $\alpha=0$ results in steering towards a constant signal with an increasing trend, whereas $\alpha=1$ only introduces sinusoidal patterns (iii). Detailed results are available in the Appendix D.
4 Results
Refer to caption Refer to caption Refer to caption
(i) MOMENT-Large (ii) Chronos-Large (iii) Moirai-1.1-R Large
Figure 6: These heatmaps illustrate pairwise similarity of layer measured using CKA. Lighter shades indicate higher similarity (dark blue $\rightarrow$ low similarity, yellow $\rightarrow$ high similarity). We see interest structures in the representations learned by these models. While MOMENT shows substantial redundancy in representations, Chronos shows a more pronounced structure of block patterns. Moirai exhibits the most distinct block patterns, suggesting clearly defined stages of representation learning. These plots offer insights into localized regions of representation utility and redundancy across the three models.
Chronos Family
(i) Tiny (ii) Mini (iii) Small (iv) Base (v) Large
Figure 7: These heatmaps demonstrate how model size influences the patterns of learned representations. Smaller models exhibit more discrete and distinct block structures, while larger models (Base and Large) display increasingly intricate and clustered patterns, reflecting more nuanced and gradual transformations across layers. Notably, the emergence of blocks-like patterns in the Large model appears unpredictable from patterns observed in smaller models.
Analyzing representations offers interesting insights.
Our analysis of model representations demonstrates that both model size and internal architecture considerably influence how representations are organized. Fig. 4) shows heatmaps which reveal that larger models, such as MOMENT-Large, Chronos-Large, and Moirai-1.1-R Large, have similar representations across specific groups of layers forming distinct and intricate block patterns, which may reflect unique stages of representation learning. More complex block patterns are observed with increasing model size, indicating that scaling may enhance the richness and organization of internal representations. However, it may also increase redundant knowledge storage through similar representations across layers, as suggested by high CKA similarity measured in block patterns. Interestingly, within model families (e.g., Chronos and Moirai), scaling does not always result in predictable heatmap changes. Larger models, like Chronos-Large and Moirai-Large, demonstrate more refined and complex transformations of representations that are not easily extrapolated from their smaller versions as shown in Fig. 14). Moreover, cross-model similarity analysis results in Fig. 8 reveal that while early layers tend to have high similarity across models of different sizes, the similarity measures among later layers diverge more notably, particularly in larger models. This divergence is especially evident in the Chronos family, where early representations are more consistent across models, but later layers become increasingly specialized as model depth increases as shown in Fig. 7.
Chronos Family
(i) Tiny vs Mini
(ii) Base vs Large
Figure 8: These heatmaps illustrate the similarity between representations learned by different layers in TSFMs of the same family but different sizes. Initial layers of these models tend to learn similar representations, while the similarity gradually decreases in the later layers.
We observed consistent improvements in memory efficiency and inference speed over their unpruned counterparts. For example, pruning only Block 3 for MOMENT-Large, resulted in a 11% decrease in estimated model size with a 5% speed up in inference. Furthermore, this pruned model had lower zero-shot imputation MAE for 5 of 9 datasets (ETTh2, ETTm1, ETTm2, Exchange, and Weather) as shown in Tab. 1. Chronos-Large results for zero-shot experiments are reported in Tab. 2. Detailed results on memory usage and speed improvements can be found in Tab. 5. While pruning consistently improved memory efficiency and inference speed compared to unpruned counterparts, performance varied across pruning methods and datasets, with some methods exhibiting considerable degradation. In addition to zero-shot experiments, we conducted experiments where models were fine-tuned post-pruning. For this, we applied MOMENT-Large for forecasting to compare a vanilla (unpruned) model to one with all block redundancies pruned, evaluating the impact of the most aggressive pruning approach. Fine-tuning results in Table 9 show that, notably, the pruned model performed nearly as well as the original, underscoring the potential of our block-wise pruning approach to maintain performance while reducing model complexity. Complete finetuning results are provided in Table 4 in Appendix D.
TSFMs learn Intuitive Linear Concepts.
Our concept localization results in Fig. 4 show that certain concepts represented by MOMENT-Large are linearly separable and that this separability is not consistent but rather emerges at specific layers in the model. We also found intuitive differences in the locations where these concepts are learned. We observed that certain concepts, such as distinguishing between constant and sinusoidal patterns, require careful examination of the entire time series. In contrast, differentiating between increasing and decreasing trends can be achieved by focusing on the initial and final patches. However, we did not identify specific locations where models learn to distinguish between time series of different amplitudes. This may be attributed to the normalization of input time series, a common practice in many TSFMs, including MOMENT.
@@ 4 Results
Forecasting Horizon
Dataset Pruning 96 192 336 720
Exchange Vanilla 0.109 0.215 0.417 1.003
All Pruned 0.113 0.218 0.394 1.066
ETTh1 Vanilla 0.385 0.411 0.423 0.443
All Pruned 0.388 0.414 0.424 0.460
ETTh2 Vanilla 0.287 0.350 0.370 0.404
All Pruned 0.296 0.356 0.382 0.404
ETTm1 Vanilla 0.290 0.330 0.352 0.409
All Pruned 0.29 0.326 0.354 0.414
ETTm2 Vanilla 0.171 0.231 0.287 0.372
All Pruned 0.173 0.236 0.294 0.372
ILI Vanilla 3.260 3.516 3.828 3.989
All Pruned 2.981 3.209 3.479 3.602
Weather Vanilla 0.153 0.197 0.246 0.316
All Pruned 0.152 0.198 0.247 0.317
Figure 9: We fine-tune the vanilla and pruned variants on MOMENT on widely used long-horizon forecasting datasets (Zhou et al., 2021) and measure MSE. We found that the pruned model performed on par with the original model, underscoring the potential of our block-wise pruning approach. In this particular pruning setup, we reduced the memory consumption by more than half compared to the vanilla model and improved inference time per sample by $\approx$1 ms.
Interestingly, the method of obtaining the steering matrix, either by computing the mean or median across embedding concept classes, has no notable effect on the steered output as shown in Fig. 12. However, applying concept steering interventions across all tokens is necessary to achieve the intended steered concept output compared to applying concept steering interventions to a single token. Moreover, the $\alpha$ parameter can have considerable effect on steered output. For Chronos, steering required tuning the parameter $\alpha\approx 0.1$ for effective performance, whereas MOMENT maintained effective steering with $\alpha=1$.
@@ 5 Discussion
We also explored ways to influence model predictions along conceptual directions using steering matrices. Concept steering has many practical applications, including the ability to correct prediction errors in pre-trained models that may arise from out-of-distribution inference, reflecting the effects of exogenous factors on model predictions or inducing prior knowledge not fully captured during training. Additionally, steering provides a method to reduce computational costs by minimizing the need for fine-tuning. It can also be used to introduce inductive biases from various domains into the model. This is particularly valuable for pre-trained models that may lack exposure to specific concepts due to limited or restricted training data. Such inductive biases can be useful in domains like healthcare, where, e.g., knowledge of the excitatory or inhibitory effects of treatments can guide pre-trained model predictions about whether a patient’s vital signs should increase or decrease in response to specific interventions. Moreover, concept steering can be used for data generation, improving data augmentation techniques and creating more diverse datasets to imxprove TSFM performance across various tasks. Prior work has already shown that data augmentation techniques can improve TSFM model generalization by enhancing model robustness and exposing it to a wider variety of patterns (Ansari et al., 2024).
@@
This paper provides insights into how a few basic patterns are linearly represented in time series foundation models. Future work must evaluate whether time series foundation models can learn more complex patterns present in real-world time series and whether steering matrices estimated using synthetic data can be used to steer predictions of out-of-distribution, real-world time series. While our methods are broadly applicable to different transformer-based foundation models, future research should explore other architectures such as state space models (Gu and Dao, 2023) and stacked multi-layer perceptrons (Ekambaram et al., 2024). Moreover, future studies should evaluate whether our findings hold for other time series foundation models and tasks such as anomaly detection, and classification. Beyond time series, we hope that our work inspires the use of synthetic data to steer large language and vision models as well.
Ethics Statement
@@ Appendix A Synthetic Data Generation
    $\alpha\sim U(\mathrm{slope}_{\min},\mathrm{slope}_{\max}),\quad\beta\sim U(\mathrm{intercept}_{\min},\mathrm{intercept}_{\max}).$
@@
    $y(t)=A\sin\left(\frac{2\pi t}{P}\right)+\alpha t+\beta,$
@@
    $A\sim U(\mathrm{amplitude}_{\min},\mathrm{amplitude}_{\max}),\quad P\sim U(\mathrm{period}_{\min},\mathrm{period}_{\max}).$
@@
    $A\sim U(0,0),\quad P\sim U(0,0),\quad\alpha\sim U(0,0),\quad\beta\sim U(-30,30).$
@@
    $A\sim U(0,0),\quad P\sim U(0,0),\quad\alpha\sim U(0.5,1),\quad\beta\sim U(-30,30).$
@@
    $A\sim U(0,0),\quad P\sim U(0,0),\quad\alpha\sim U(-1,-0.5),\quad\beta\sim U(-30,30).$
@@
    $A\sim U(50,50),\quad P\sim U(128,128),\quad\alpha\sim U(0,0),\quad\beta\sim U(-30,30).$
@@
    $A\sim U(50,50),\quad P\sim U(128,128),\quad\alpha\sim U(0.5,1),\quad\beta\sim U(-30,30).$
@@
    $A\sim U(50,50),\quad P\sim U(128,128),\quad\alpha\sim U(-1,-0.5),\quad\beta\sim U(-30,30).$
Constant Pattern ($y(t)=\alpha t+\beta$)
Refer to caption Refer to caption Refer to caption
(i) No trend (ii) Trend (vary $t$)
Sinusoidal Pattern ($y(t)=A\sin(\frac{2\pi t}{P})+\alpha t+\beta$)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(iii) Amplitude (vary $A$) (iv) Periodicity (vary $P$) (v) Trend (vary $\alpha$)
Figure 10: Samples of synthetic time series used in our experiments. We use two base patterns (constant and sinusoidal). To generate synthetic datasets, we vary the periodicity ($f$), amplitude ($a$), intercept ($c$), and linear trend ($t$).

### v3

@@ 3.3 Concept-Informed Predictions via Model Steering
Once we have identified that a feature is linearly represented in the latent space of the $\mathcal{M}$, we can use steering interventions to manipulate the latent space and generate time series that reflect intended concepts. For instance, to introduce periodicity to a constant time series, we can utilize a steering matrix $\mathbf{S}$, as illustrated in Figure 2. By strategically intervening in $\mathcal{M}$ using this steering matrix, we can bias its outputs towards predicting periodic time series. To construct a steering matrix, we first derive steering vectors $\mathbf{S}_{i}\in\mathbb{R}^{N\times D}$, for each layer $i$. These vectors represent the change that activations in layer $i$ must undergo such that $\mathcal{M}$ produces periodic outputs. $\mathbf{S}_{i}$ is simply the difference between the median activation matrix of the constant time series $\mathbf{M}_{i_{c}}$, from that of sinusoids $\mathbf{M}_{i_{s}}$. We stack these vectors for all layers to derive the steering matrix. This matrix allows us to simultaneously intervene across multiple tokens and layers during inference, which we found to be more effective than single-token interventions. During inference, to steer model predictions, at each layer $i$, we update its hidden representation as follows: $\mathbf{h}_{i}\leftarrow\mathbf{h}_{i}+\lambda\mathbf{S}_{i},$ where $\lambda\in\mathbb{R}$ is a scalar that controls the strength of the intervention.
@@
We explore two alternative modalities of intervention: (1) deriving steering vectors using the mean of hidden activations rather than the median, and (2) steering a single token versus all tokens throughout the model. While our methods are applicable to a wide range of transformer-based foundation models, we focus on two prominent TSFM families for brevity: MOMENT (Goswami et al., 2024) and Chronos (Ansari et al., 2024). Both these models are fully open-source, come in different sizes, yet have fundamentally different design choices. For example, Chronos is based on encoder-decoder transformer (in our work we will investigate encoder part) model which takes discretized time series as input, whereas MOMENT is a multi-task, encoder-only model which takes continuous time series patches as input. Since only the Large variant of MOMENT is publicly available at the time of writing this paper, we supplement our representation analysis results with Moirai (Woo et al., 2024), another popular TSFM which comes in different sizes. More information on hyper-parameters can be found in Appendix E.
4 Results
Analyzing representations offers interesting insights.
Our analysis of model representations demonstrates that both model size and internal architecture considerably influence how representations are organized. Fig. 7 shows heatmaps which reveal that larger models, such as MOMENT-Large, Chronos-Large, and Moirai-1.1-R Large, have similar representations across specific groups of layers forming distinct and intricate block patterns, which may reflect unique stages of representation learning. More complex block patterns are observed with increasing model size, indicating that scaling may enhance the richness and organization of internal representations. However, it may also increase redundant knowledge storage through similar representations across layers, as suggested by high CKA similarity measured in block patterns. Interestingly, within model families (e.g., Chronos and Moirai), scaling does not always result in predictable heatmap changes. Larger models, like Chronos-Large and Moirai-Large, demonstrate more refined and complex transformations of representations that are not easily extrapolated from their smaller versions as shown in Fig. 15). Moreover, cross-model similarity analysis results in Fig. 4 reveal that while early layers tend to have high similarity across models of different sizes, the similarity measures among later layers diverge more notably, particularly in larger models. This divergence is especially evident in the Chronos family, where early representations are more consistent across models, but later layers become increasingly specialized as model depth increases as shown in Fig. 6.
We observed consistent improvements in memory efficiency and inference speed over their unpruned counterparts. For example, pruning only Block 3 for MOMENT-Large, resulted in a 11% decrease in estimated model size with a 5% speed up in inference. Furthermore, this pruned model had lower zero-shot imputation MAE for 5 of 9 datasets (ETTh2, ETTm1, ETTm2, Exchange, and Weather) as shown in Tab. 3. Chronos-Large results for zero-shot experiments are reported in Tab. 4. Detailed results on memory usage and speed improvements can be found in Tab. 7. While pruning consistently improved memory efficiency and inference speed compared to unpruned counterparts, performance varied across pruning methods and datasets, with some methods exhibiting considerable degradation. In addition to zero-shot experiments, we conducted experiments where models were fine-tuned post-pruning. For this, we applied MOMENT-Large for forecasting to compare a vanilla (unpruned) model to one with all block redundancies pruned, evaluating the impact of the most aggressive pruning approach. Fine-tuning results in Table 1 show that, notably, the pruned model performed nearly as well as the original, underscoring the potential of our block-wise pruning approach to maintain performance while reducing model complexity. Complete finetuning results are provided in Table 6 in Appendix F.
Chronos Family
Refer to caption Refer to caption
(i) Tiny vs Mini (ii) Base vs Large
Figure 4: Similarity between representations learned by different layers in TSFMs of the same family but different sizes. Initial layers tend to learn similar representations, while the similarity gradually decreases in the later layers.
Refer to caption Refer to caption Refer to caption
(i) MOMENT (ii) Chronos (iii) Moirai
Figure 5: Pairwise similarity of layer measured using CKA for large variants of TSFMs (dark blue $\rightarrow$ low similarity, yellow $\rightarrow$ high similarity). All TSFMs learn redundant representations with manifesting as block-like structures.
ETTh1 ETTh2 ETTm1 ETTm2 ILI Weather
Vanilla 0.385 0.287 0.290 0.171 3.260 0.153
Pruned 0.388 0.296 0.290 0.173 2.981 0.152
Table 1: MSE of fine-tuned vanilla and pruned MOMENT variants on long-horizon forecasting datasets (Zhou et al., 2021). Pruned models perform on par with the original, whilereducing memory consumption by $>50\%$ and improving inference time per sample by $\approx$1 ms. More comprehensive results are available in the appendix in Tables 6 and 7.
Chronos Family
(i) Tiny (ii) Mini (iii) Small (iv) Base (v) Large
Figure 6: How does model size influence the patterns of learned representations? The emergence of blocks-like patterns in the Large model appears unpredictable from patterns observed in smaller models. More available in the appendix (Fig. 16).
Refer to caption Refer to caption Refer to caption
(i) Pattern (ii) Trend (iii) Periodicity
Figure 7: Progression of linear separability of concepts at the patch level (y-axis) across different layers (x-axis). Linear separability is measured using the Linear Discriminant Ratio (LDR), derived from model embedding statistics for each predicted class: (i) constant vs. sinusoidal patterns, (ii) increasing vs. decreasing trends, and (iii) high vs. low periodicity. Color gradient indicates separability, with dark blue representing low LDR and yellow indicating high LDR. Certain concepts captured by MOMENT-Large exhibit linear separability, but this separability is not uniform across layers; instead, it emerges at specific points in the model. More available in the appendix (Fig. 18).
TSFMs learn intuitive linear concepts.
Our concept localization results in Fig. 7 show that certain concepts represented by MOMENT-Large are linearly separable and that this separability is not consistent but rather emerges at specific layers in the model. We also found intuitive differences in the locations where these concepts are learned. We observed that certain concepts, such as distinguishing between constant and sinusoidal patterns, require examination of the entire time series. In contrast, differentiating between increasing and decreasing trends can be achieved by focusing on the initial and final patches. However, we did not identify specific locations where models learn to distinguish between time series of different amplitudes. This may be attributed to the normalization of input time series, a common practice in many TSFMs, including MOMENT.
@@ 4 Results
(i) Trend, $\beta=1.0$ (ii) Trend & periodicity, $\beta=0.5$ (iii) Periodicity, $\beta=0$
@@
Figure 8: Visualization of the MOMENT’s reconstruction and the Chronos’s forecasting predictions (bottom), with concept steering applied in the latent space (blue) and the baseline without concept steering (orange). Concept steering effectively transforms concepts in the latent space, resulting in model predictions that align with the intended concepts introduced. Steering results are illustrated for the following experiments: steering a constant signal input to produce (i) a sinusoidal output, (ii) a constant signal with an increasing(slope $>$ 0), and (iii) decreasing trend (slope $<$ 0). For compositional steering experiments, $\alpha$ controls the strength of sinusoidal and increasing trend concepts. When $\beta=0.5$, models are steered towards a combination of increasing trend and sinusoidal pattern. $\beta=0$ results in steering towards a constant signal with an increasing trend, whereas $\beta=1$ only introduces sinusoidal patterns (iii). Detailed results are available in the Appendix F.
@@
Interestingly, the method of obtaining the steering matrix, either by computing the mean or median across embedding concept classes, has no notable effect on the steered output as shown in Fig. 11. However, applying concept steering interventions across all tokens is necessary to achieve the intended steered concept output compared to applying concept steering interventions to a single token. Moreover, the $\lambda$ parameter can have considerable effect on steered output. For Chronos, steering required tuning the parameter $\lambda\approx 0.1$ for effective performance, whereas MOMENT maintained effective steering with $\lambda=1$.
@@ 5 Discussion
We also explored ways to influence model predictions along conceptual directions using steering matrices. Concept steering has many practical applications, including the ability to correct prediction errors in pre-trained models that may arise from out-of-distribution inference, reflecting the effects of exogenous factors on model predictions or inducing prior knowledge not fully captured during training. Additionally, steering provides a method to reduce computational costs by minimizing the need for fine-tuning. It can also be used to introduce inductive biases from various domains into the model. This is particularly valuable for pre-trained models that may lack exposure to specific concepts due to limited or restricted training data. Such inductive biases can be useful in domains like healthcare, where, e.g., knowledge of the excitatory or inhibitory effects of treatments can guide pre-trained model predictions about whether a patient’s vital signs should increase or decrease in response to specific interventions. Moreover, concept steering can be used for data generation, improving data augmentation techniques and creating more diverse datasets to improve TSFM performance across various tasks. Prior work has already shown that data augmentation techniques can improve TSFM generalization by enhancing model robustness and exposing it to a wider variety of patterns (Ansari et al., 2024).
@@
This paper provides insights into how a few basic patterns are linearly represented in time series foundation models. Future work must evaluate whether time series foundation models can learn more complex patterns present in real-world time series and whether steering matrices estimated using synthetic data can be used to steer predictions of out-of-distribution, real-world time series. While our methods are broadly applicable to different transformer-based foundation models, future research should explore other architectures such as state space models (Gu & Dao, 2023) and stacked multi-layer perceptrons (Ekambaram et al., 2024). Moreover, future studies should evaluate whether our findings hold for other time series foundation models and tasks such as anomaly detection, and classification. Beyond time series, we hope that our work inspires the use of synthetic data to steer large language and vision models as well.
Acknowledgment
This work was partially supported by NSF (awards 2406231 and 2427948), NIH (awards R01NS124642, R01DK131586, and R01NR013912), and the US Army (W911NF-20-D0002).
Impact Statement
@@ Appendix A Synthetic Data Generation
Constant Pattern ($y(t)=mt+b$)
Refer to caption Refer to caption Refer to caption
(i) No trend (ii) Trend (vary $t$)
Sinusoidal Pattern ($y(t)=a\sin(\frac{2\pi t}{f})+mt+b$)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(iii) Amplitude (vary $a$) (iv) Periodicity (vary $f$) (v) Trend (vary $m$)
Figure 9: Samples of synthetic time series used in our experiments. We use two base patterns (constant and sinusoidal). To generate synthetic datasets, we vary the periodicity $f$, amplitude $a$, intercept $b$, and linear trend $m$.
@@
    $m\sim U(\mathrm{slope}_{\min},\mathrm{slope}_{\max}),$
    $b\sim U(\mathrm{intercept}_{\min},\mathrm{intercept}_{\max}).$
@@
    $y(t)=a\sin\left(\frac{2\pi t}{f}\right)+mt+b,$
@@
    $a\sim U(\mathrm{amplitude}_{\min},\mathrm{amplitude}_{\max}),$
    $f\sim U(\mathrm{period}_{\min},\mathrm{period}_{\max}).$
@@
    $a\sim U(0,0),\quad f\sim U(0,0),\quad m\sim U(0,0),\quad b\sim U(-30,30).$
@@
    $a\sim U(0,0),\quad f\sim U(0,0),\quad m\sim U(0.5,1),\quad b\sim U(-30,30).$
@@
    $a\sim U(0,0),\quad f\sim U(0,0),\quad m\sim U(-1,-0.5),\quad b\sim U(-30,30).$
@@
    $a\sim U(50,50),\quad f\sim U(128,128),\quad m\sim U(0,0),\quad b\sim U(-30,30).$
@@
    $a\sim U(50,50),\quad f\sim U(128,128),\quad m\sim U(0.5,1),\quad b\sim U(-30,30).$
@@
    $a\sim U(50,50),\quad f\sim U(128,128),\quad m\sim U(-1,-0.5),\quad b\sim U(-30,30).$